Find the measures of the angles of each triangle:
The ratio of the measures of 3 angles is 2:5:3

Answer by Alan3354(69443)

2 + 5 + 3 = 10
The sum is 180
180/10 = 18
2*18 = 36º
5*18 = 90º
3*18 = 54º